6 years ago I was 11 times as old as my daughter. Now I am 5 times as old as she is. How old am I?

Answer by ptaylor(2198)

6 years ago I was 11 times as old as my daughter. Now I am 5 times as old as she is. How old am I?
Let x=daughter's current age-----x-6=daughter's age 6 yrs ago
Then 5x=your current age----5x-6=your age 6 yrs ago
Soooo:
5x-6=11(x-6)
5x-6=11x-66
11x-5x=66-6
6x=60
x=10 daughter's age now
5x=5*10=50-----your current age
CK
50-6=11(10-6)
44=44
Hope this helps---ptaylor